Police report that 90% of drivers stopped on suspicion of drunk driving are given a
breath test, 11% are given a blood test, and 8% are given both.
(a) In this context, define two events A and B.
(b) Write the given information in probability notation.
(c) Explain, in this context, the meaning of P(A ∩ B).
(d) Are A and B disjoint events?
(e) Are A and B independent events?
1
Expert's answer
2019-11-19T09:25:04-0500

Discrete Mathematics



(a).


Let the Event A = Drivers given the breath test for drunk driving


Event B = Drivers given the blood test for drunk driving


(b)


Probability of event A = P(A) = 90% = "\frac {90} {100} = 0.9"


Probability of event B = P(B) = 11% = "\frac {11}{100} = 0.11"


P(A ∩ B)= 8% = 0.08


(c).

P(A ∩ B)= The probability that the drivers stopped on suspicion of drunk driving are both given both breath test

and blood test


(d)


If P(A ∩ B)= 0, then A and B are disjoint events .


Here, P(A ∩ B)= 0.08


So, A and B are not disjoint events.


(e)


If "P(A ∩ B)= P(A) \times P(B)," then A and B are independent events.


P(A ∩ B)= 0.08


"P(A) \times P(B) = 0.9 \times 0.11 = 0.099"


"P(A ∩ B) \ne P(A) \times P(B)"


Answer: So, A and B are not independent events.